1. Find a formula for the function whose graph consists of the line segment from the point (−4,3) to the point (−2,0) and the lower half of the circle centered at the origin with radius 2.
1
Expert's answer
2021-09-21T05:00:10-0400

To determine:

An expression function for the graph which satisfies the given condition:

Given:

The graph has a line segment connecting (-4,3) and (-2,0) and it consists of a top half of the circle with center (0,0) and radius 2.

Calculation:

Find the slop of the line segment joining the points (-4,3) and (-2,0) as follows:

"m={y_2-y_1 \above{2pt} x_2-x_1}"


"m={0-3 \above{2pt} -2-(-4)}"

"m={-3 \above{2pt} 2}"

Thus, the slope of the line segment is "m=-\frac{3}{2}" .

Find the y-intercept of the line segment joining the points (-4,3) and (-2,0) as follows

"y=mx+c"

"0=(-\frac{3}{2})(-2)+c" 

"0=3+c"

"c=-3"

Thus , y intercept is c=-3.

The equation of the line segment is "y=-\frac{3}{2}x-3" for "-4\leqslant x<-2" .

Find the equation of the circle:

The equation of the circle of radius 2 centered on the origin is "x^2+y^2=4"

We can then solve for y

"x^2+y^2=4\iff" "y^2=4-x^2\iff" "y=\sqrt{4-x^2}"

To only include the top half of the circle we only take the positive root .

The equation of the circle is "\sqrt{x^2-4}" for -22

"f(x)=\begin{cases}
 -\frac{3}{2}x-3 &\text{ on} [-4,-2) \\
 \sqrt{4-x^2} &\text{on} [-2,2]
\end{cases}"